Session: Nov 2001 (session 3)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Huang Xinyong at Fri Nov 23 02:22:59 MET 2001) Success
The Tsys of the RCP is higher than LCP.

Urumqi (Liu Xiang at Fri Nov 23 04:04:29 MET 2001) Success
H-maser to GPS 2.212us, 2.199us, formatter to GPS 1.987us, 2.19us.

Effelsberg (A. Kraus at Fri Nov 23 09:04:37 MET 2001) Success

Torun (Kaz Borkowski at Fri Nov 23 09:13:07 MET 2001) Success
No known problems.

Onsala (John Conway at Fri Nov 23 11:08:21 MET 2001) Minor failures
No problems except the logs are apparently lost. The expt was started, the tape needed to be remounted, the log was deleted using rm in the log directory (since it contained rubbish) and then the schedule was restarted in the FS but apparently no new log file was opened. Ivestigating to see if it exists anywhere on the FS.

Jodrell_Bank (Alastair Gunn at Sun Nov 25 20:59:46 MET 2001) Success
No known problems. Faulty Tsys from VC06, but data okay.

\(Other\) (G.Maccaferri at Wed Nov 28 10:18:49 MET 2001) Success

Medicina (G.Maccaferri at Wed Nov 28 10:19:12 MET 2001) Success